inlet to Cranberry juice outlet
media%2F184%2F184fd1af-7652-4227-a2c4-34
Answer #1

The beginning of the straw is below the surface of the cranberry juice. The end of the straw is a point above the surface of the orange juice. Let h_o be the height differential between the surface of cranberry juice and the end point of the straw located above the orange juice glass."

Anyhow, with this reading, it will be v = sqrt (2*g*h_o)

The easy way this is derived is that the Kinetic Energy at exit is equal to the potential energy of the drop in elevation. Basically 1/2 mv^2=mgh or v=sqrt (2gh)